OpenStudy (anonymous):

Integrate:

OpenStudy (anonymous):

\[\int\limits_{d}^{0}\frac{\sqrt{x}}{\sqrt{d-x}}dx\]

OpenStudy (dumbcow):

the limits cant be right...function is undefined for x<0

OpenStudy (zzr0ck3r):

good call

OpenStudy (anonymous):

d>0

OpenStudy (dumbcow):

domain is 0<x<d

OpenStudy (anonymous):

i do not think anything is wrong. The book says to substitute x=u\(^{2}\).

OpenStudy (dumbcow):

ok then dx = 2u du \[2 \int\limits \frac{\sqrt{u^{2}}*u}{\sqrt{d-u^{2}}} du\]

OpenStudy (dumbcow):

then make sub ... u = sqrt(d)*sin w du = sqrt(d)*cos w dw \[2\int\limits \frac{d \sin^{2} w}{\sqrt{d} \sqrt{1-\sin^{2} w}}\sqrt{d} \cos w dw\]

OpenStudy (dumbcow):

\[=2d \int\limits \sin^{2} w dw\]

OpenStudy (dumbcow):

do you know integral of sin^2 ?

OpenStudy (anonymous):

i think i got it. Thanks!

OpenStudy (dumbcow):

yw

OpenStudy (anonymous):

Got the answer! Thanks again.